A ray of light is minimally deviated when incident at an angle 30 degrees on one face of a glass prism of refractive index 1.5. The minimum angle of deviation is?
1
Expert's answer
2020-07-29T09:49:39-0400

Solution;

"n = \frac{sin i}{sin r}"


"sin r = \frac{30^o}{n}"


"sin r = \frac{0.5}{1.5}"

Then, "sin r = 0.3333"

solving you find "r=19.5\degree"

Therefore deviation is; "30\degree-19.5=10.5\degree"